Find all points on the graph of the equation x2-y2=2x+4y where the tangent line is horizontal. Does the graph have any vertical asymptote? Sketch the graph
1
Expert's answer
2020-07-07T19:39:07-0400

x2-y2=2x+4y

x2-y2(x)-2x-4y(x)=0

=>

(x2-y2(x)-2x-4y(x))'=0

2x-2y(x)y'(x)-2-4y'(x)=0

The condition y'(x)=0 is true in points where the tangent line is horizontal.

If y'(x)=0 then 2x-2=0 => x=1

12-y2-2*1-4y=0 =>

y2+4y+1=0

y1=-2-31/2, y2=-2+31/2

(1;-2-31/2), (1;-2+31/2) are points where the tangent line is horizontal.







(x-1)2-(y+2)2=-3

x-1=y+2 =>y=x-3

x-1=-(y+2) => y=-x-1

Asymptotes :

y=x-3, y=-x-1

So, the graph does not have any vertical asymptote.

Answer:

(1;-2-31/2), (1;-2+31/2) are points where the tangent line is horizontal;

The graph does not have any vertical asymptote